# ArcMemo: Abstract Reasoning Composition with Lifelong LLM Memory

### v2

@@ Abstract
While inference-time scaling enables LLMs to carry out increasingly long and capable reasoning traces, the patterns and insights uncovered during these traces are immediately discarded once the context window is reset for a new query. External memory is a natural way to persist these discoveries, and recent work has shown clear benefits for reasoning-intensive tasks. We see an opportunity to make such memories more broadly reusable and scalable by moving beyond instance-based memory entries (e.g. exact query/response pairs, or summaries tightly coupled with the original problem context) toward concept-level memory: reusable, modular abstractions distilled from solution traces and stored in natural language. For future queries, relevant concepts are selectively retrieved and integrated into the prompt, enabling test-time continual learning without weight updates. Our design introduces new strategies for abstracting takeaways from rollouts and retrieving entries for new queries, promoting reuse and allowing memory to expand with additional experiences. We evaluate on ARC-AGI, a benchmark that stresses compositional generalization and abstract reasoning, making it a natural fit for concept memory. Our method yields a 7.5% relative gain over a strong no-memory baseline with performance continuing to scale with inference compute. We find abstract concepts to be the most consistent memory design, outscoring the baseline at all tested inference compute scales. Moreover, dynamically updating memory during test-time outperforms fixed settings, supporting the hypothesis that accumulating and abstracting patterns enables further solutions in a form of self-improvement  [^1] Code available at https://github.com/matt-seb-ho/arc_memo.
@@ 1 Introduction
Large language models (LLMs) have made substantial strides in reasoning-intensive tasks with long-form reasoning. However, a notable opportunity lies in the fact that LLM systems are fixed after deployment: new patterns and strategies uncovered during deep reasoning are not yet carried forward once the context is cleared. This contrasts with human approaches to solving complex, compositional reasoning problems, which involve building on prior insights, abstracting patterns, and composing them in new contexts. Augmenting LLMs with memory offers a solution to carrying forward and building on their discoveries.
While external augmentation was initially designed for factuality in knowledge-intensive tasks (e.g. Lewis et al. (2021) with RAG, Zhang et al. (2019) with Knowledge Graphs, inter alia), recent work has begun developing memory approaches for reasoning-intensive tasks, storing questions, findings, and mistakes in memory instead of simple facts (Yang et al., 2024). These memories tend to be specific to the problem/experience it was originally derived from (we term these “instance-level” concepts, see Figure 1). Although effective for closely related problems, they have diminished utility for problems superficially different from prior experiences. Other approaches such as Suzgun et al. (2025) begin addressing the “instance-level” issue by maintaining an evolving summary of previous experience to great effect, but the lack of structure and modularity makes scaling with more experiences challenging. Without selective retrieval, the size of memory is limited and the problem solving model has the additional burden of picking out the relevant ideas from all past experiences.
@@
We evaluate on ARC-AGI, where simple pixel-grid operations compose into a vast task space, and solving tasks requires compositional reasoning rather than memorizing individual patterns, making it a natural testbed for ArcMemo. On ARC-AGI-1, our method improves the official score from 55.17 to 59.33 (+7.5% relative gain over a strong no-memory baseline) and is the only memory design we find to outperform the baseline at all inference scales. Our experiments confirm that continually updating memory at evaluation time yields a score improvement that emerges with inference scale via retries: memory updates triggered from a previous pass over the test set may enable new solves in a subsequent pass. Finally, we observe that selecting a subset of memories to include for a particular problem improves performance and reduces token cost, indicating the selection operation is essential beyond allowing memory to continually grow.
@@ 2 Related Work
Our approach draws on several threads of research in augmenting LLMs with memory but contrasts in (1) target applications (reasoning-intensive vs. knowledge-intensive tasks), (2) the underlying memory modality (text vs. continuous vectors/embeddings), and (3) emphasis on abstraction and modularity for reasoning. We discuss related lines (1) and (2) as well as other related work in Appendix D and emphasize reasoning and abstraction here.
@@
More recent work has shifted toward reasoning-centric uses of memory, particularly test-time learning. Think-in-Memory records intermediate reasoning steps as structured triples, storing them in a locality-sensitive hash table to enable reuse in multi-step reasoning tasks (Liu et al., 2023). Buffer of Thoughts stores reusable reasoning templates, retrieved by embedding similarity and instantiated for new problems (Yang et al., 2024). It categorizes and summarizes solution attempts into templated insights, which are then added to memory. In contrast, Dynamic Cheatsheet maintains a unified buffer that is adapted continuously. With each problem query, the LLM updates this memory blob by rewriting the entire buffer (Suzgun et al., 2025). Rather than retrieving specific entries, the entire cheatsheet is appended to the prompt, functioning as a persistent cache of problem-solving strategies.
@@
Our work is primarily evaluated against ARC-AGI as a benchmark that simulates complex reasoning on frontier tasks without requiring frontier-level knowledge. Various techniques and findings have been developed against this benchmark. Li et al. (2024) demonstrate the efficacy of synthetic data and the complementary approaches of (i) learning to infer the underlying transformation via program synthesis and (ii) using test time weight adaptation to directly learn the transformation function in the neural net. Akyürek et al. (2025) further investigates the test time adaptation method, producing various insights, including that low rank adapters are poorly suited to retaining useful experience across puzzles. Our program synthesis approach can be viewed as a memory-augmented version of Wang et al. (2024)’s hypothesis search or Qiu et al. (2024)’s hypothesis refinement with execution feedback based retry. Other recent approaches to ARC-AGI include brain-inspired architectural innovations such as Wang et al. (2025)’s hierarchical reasoning model achieving impressive performance through a recurrent architecture that iteratively refines its output until a halt signal is predicted together with data augmentation techniques. Finally, the relative ease of generating new ARC-AGI tasks compared to solving them is exploited in Pourcel et al. (2025)’s self-improving system that relabels incorrect programs as the correct program for a different puzzle and trains a model on its generations this way.
@@ 3.1 Problem Statement
Input: Dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{n}$ (labels $y_{i}$ optional), External memory $M$, Operations MemRead, MemWrite, and GetFeedback, Update interval $k$,
@@
Algorithm 1 Inference with External Memory
Figure 2: Method Diagram. Implementing a memory system requires defining (1) what is stored (2) how memory is updated (3) how memory is used for new queries. The key novelties in this work is emphasizing abstraction and modularity and the corresponding design changes. In particular we highlight that parameterization (with higher-order functions allowed and encouraged) promotes abstraction and typed interface definitions supports modularity by showing which concepts can be combined. Since these memory entries are more abstract, they also require more inference to map against new, concrete situations– whether by aligning input against the memory format, or by instructing a reasoning model to explore.
@@
The main requirement for our memory system is that some feedback is available at test time (e.g. via test cases, or self-reflection). This condition is necessary as retaining patterns from an incorrect or otherwise flawed trace would serve to carry mistakes forward. In other words, some signal is needed to discriminate correct from incorrect traces to ensure only productive ideas and patterns are added to memory. Incorrect traces still likely contain vital signal, but error identification/credit assignment remains an open challenge we leave to future efforts. Various real-world tasks may satisfy this requirement, as tasks are often specified with either examples showing desired behaviour or evaluation criteria defining what outcomes are positive and negative. To give some examples, code completion tools have criteria like compilation success and test pass rate, and medical diagnosis tools have evaluation criteria defined as patient outcomes (e.g., survival rates). In the case of our evaluation benchmark ARC-AGI, puzzles are explicitly presented as a set of input-output examples which for our purposes, can serve as automatically verifiable test cases.
@@ 3.2 Memory Format and Organization
Specifying an individual memory entry’s format involves selecting fields useful for retrieval and downstream problem solving. The organization of memory entries on a collective level is another design surface we leave for future exploration; we use a flat collection of entries for simplicity.
Figure 3: Open-Ended (OE) vs. Program Synth (PS) Concept Examples. A example of concepts from puzzle 9af7a82c abstracted into each concept format. OE defers to the model, while PS imposes structure to encourage abstraction/modularity. Higher order behaviour is demonstrated with the “sort objects” concept taking a Callable parameter that specifies specific variations.
@@
In early iterations of automatic concept summarization from solution traces, we observed outputs to occasionally be overly specific. Consider Figure 3’s OE example that bundles ideas of counting, drawing columns, and sorting into a single entry. This overspecification inherits the same limitations as instance-level concepts: friction when matching a composite idea to new scenarios and redundancy from composites ideas sharing components.
We designed a more structured alternative to the OE format to address these issues and more strongly enforce the ideals of abstraction and modularity. In particular, we took inspiration from software engineering and functional programming which respectively have existing solutions in modular design for code reuse and higher level functions for composition. From this analogy, our Program Synthesis (PS) format frames concepts as types, structures, and routines. The primary feature is that each of these concepts are parameterized. Parameterization allows similar concepts to be represented compactly with variations abstracted into parameters. Moreover, type annotations for inputs (parameters) and outputs (return types) encourage modularity and composition: the typed interfaces suggest what ideas fit together and how. Finally, by allowing parameters to be routines themselves (introducing higher-order functions), we encourage further generalization by enabling abstraction over specialized logic and routines rather than just values. Higher-level patterns can be recognized across instances and stored with routine arguments defining the different lower-level logic. A subtle benefit of this more structured formulation is that memory representation can be simply compressed omitting certain fields (see subsection C.1 for a complete list of PS concept fields).
@@ 3.3 Memory Write: Concept Abstraction
The main purpose of memory is to persist discoveries and reflections from prior experiences. Converting experiences (rollouts) to memory updates is then the most crucial operation for achieving persistence.
@@
Abstracting suggestions from a solution trace is straightforward–simply query a model to reflect on the solution trace and summarize certain general ideas that may be reused for future puzzles. To synthesize specific conditions or situational cues to pair with suggestions, ideally we can directly refer to the original problem-solving process as a series of deductions that derive new facts or explore ideas from existing facts, and link each suggestion (idea) with the prior observations that inferred this suggestion. However, intermediate inferences may be implicit, or difficult to extract from a trace, or in the case of certain commercial models, completely hidden. To address this challenge, we generate a post-hoc derivation: an interleaved sequence of observations and thoughts/reasoning steps constructed from the input-output of the final solution. These reconstructed traces are then used to extract situation-suggestion pairs, forming structured memory entries for guiding future analyses. This minimal formulation does not consider the existing contents of memory while adding new entries for simplicity and scaling reasons. It defers handling redundancy and consolidating ideas to the concept selection and problem solving phases respectively.
@@
We find that directly converting solutions into routines often leads to minor implementation details being recorded as memories as opposed to our intended abstract concepts. We mitigate this by preprocessing solutions into pseudocode to prioritize higher level operations over low level implementation details. The main abstraction phase proceeds from the generated pseudocode, recording new concepts and revising existing concepts along with their various fields.
@@ 3.4 Memory Read: Concept Selection
Although state-of-the-art models support longer context windows, its maximum length still bounds how much information or how many memory entries can be included in a single query. Even without this hard limit, including all available memories into context is still undesirable: it can distract the model with irrelevant details and flood it with too many unlikely hypotheses to consider. A more effective strategy is to introduce a selection mechanism to only include the most relevant subset of memory entries at problem solving time. This allows the memory store to grow continually without overwhelming the model’s context window.
@@
OE-format memory entries were explicitly designed to support memory-based selection. The situation clause acts as a semantic hook to identify relevance. The key idea for OE Selection is a preprocessing step that leverages model reasoning capability to parse the abstract domain input into problem descriptions at different levels of abstraction to facilitate matching against our abstract concepts. Since ARC-AGI deals with spatial reasoning, we leverage a vision language model for this preprocessing step. We caption each puzzle using a structured prompt that separates concrete observations from speculative transformations. This converts spatially rich input into a natural language format suitable for matching against stored situations. We then query a model for the top-$k$ most relevant entries using the generated description. We explored scoring, thresholding, and cumulative similarity approaches (as in top-$p$ sampling). Observing comparable results, we selected top-$k$ cutoff for simplicity.
We consider this a light-weight approach to converting abstract domain inputs into more understandable and memory-aligned representations.
@@
Standard embedding approaches use a single forward pass per embedding, representing a System 1-style fast thinking/intuition driven understanding from the model. Intuitions may not be sufficient for difficult/frontier tasks and domains. Moreover, the connection between a higher level concept with a specific problem instance may not be immediately clear because of the abstraction.
To address these challenges, we propose reasoning-based selection– using System 2-style thinking to deliberately think and explore. In contrast to the more straightforward playbooks curated in the OE memory format, the explicitly abstract PS concepts are a set of puzzle pieces with notes on identifying if a piece is relevant (relevant cues) and on how various pieces fit together (type annotations). To select from this memory, we propose leveraging model-driven exploration instantiated by recent models’ long-form reasoning with backtracking. PS Selection instructs a reasoning model to systematically explore the problem: first identify initial concepts using relevance cue annotations, then attempt to “fill in the details” by determining values or routines to populate these initial concepts’ parameters using type annotations to identify which other concepts should be investigated.

### v3

@@ Abstract
While inference-time scaling enables LLMs to carry out increasingly long and capable reasoning traces, the patterns and insights uncovered during these traces are immediately discarded once the context window is reset for a new query. External memory is a natural way to persist these discoveries, and recent work has shown clear benefits for reasoning-intensive tasks. We see an opportunity to make such memories more broadly reusable and scalable by moving beyond instance-based memory entries (e.g. exact query/response pairs, or summaries tightly coupled with the original problem context) toward concept-level memory: reusable, modular abstractions distilled from solution traces and stored in natural language. For future queries, relevant concepts are selectively retrieved and integrated into the prompt, enabling test-time continual learning without weight updates. Our design introduces new strategies for abstracting takeaways from rollouts and retrieving entries for new queries, promoting reuse and allowing memory to expand with additional experiences. We evaluate on ARC-AGI, a benchmark that stresses compositional generalization and abstract reasoning, making it a natural fit for concept memory. Our method yields a 7.5% relative gain over a strong no-memory baseline, with performance continuing to scale with inference compute. We find abstract concepts to be the most consistent memory design, outscoring the baseline at all tested inference compute scales. Moreover, dynamically updating memory during test-time outperforms fixed settings, supporting the hypothesis that accumulating and abstracting patterns enables further solutions in a form of self-improvement.  [^1] Code available at https://github.com/matt-seb-ho/arc_memo.
@@ 1 Introduction
Large language models (LLMs) have made substantial strides in reasoning-intensive tasks with long-form reasoning. However, a notable opportunity lies in the fact that LLM systems are fixed after deployment: new patterns and strategies uncovered during deep reasoning are not yet carried forward once the context is cleared. This contrasts with human approaches to solving complex, compositional reasoning problems, which involve building on prior insights, abstracting patterns, and composing them in new contexts. Augmenting LLMs with memory offers a natural solution to retaining and building on their discoveries.
While external augmentation was initially designed for factuality in knowledge-intensive tasks (e.g. Lewis et al. (2021) with RAG, Zhang et al. (2019) with Knowledge Graphs, inter alia), recent work has begun developing memory approaches for reasoning-intensive tasks, storing questions, findings, and mistakes in memory instead of simple facts (Yang et al., 2024). These memories tend to be specific to the problem/experience it was originally derived from (we term these “instance-level” concepts, see Figure 1). Although effective for closely related problems, they have diminished utility for problems superficially different from prior experiences. Other approaches, such as Suzgun et al. (2025), begin addressing the “instance-level” issue by maintaining an evolving summary of previous experience to great effect, but the lack of structure and modularity makes scaling with more experiences challenging. Without selective retrieval, the size of memory is limited, and the problem solving model has the additional burden of picking out the relevant ideas from all past experiences.
@@
We evaluate on ARC-AGI, where simple pixel-grid operations compose into a vast task space, and solving tasks requires compositional reasoning rather than memorizing individual patterns, making it a natural testbed for ArcMemo. On ARC-AGI-1, our method improves the official score from 55.17 to 59.33 (+7.5% relative gain over a strong no-memory baseline) and is the only memory design we find to outperform the baseline at all inference scales. Our experiments confirm that continually updating memory at evaluation time yields a score improvement that emerges with inference scale via retries: memory updates triggered from a previous pass over the test set may enable new solves in a subsequent pass. Finally, we observe that selecting a subset of memories to include for a particular problem improves performance and reduces token cost, indicating the selection operation is essential beyond allowing memory to grow continually.
@@ 2 Related Work
Our approach draws on several threads of research in augmenting LLMs with memory but contrasts in (1) target applications (reasoning-intensive vs. knowledge-intensive tasks), (2) the underlying memory modality (text vs. continuous vectors/embeddings), and (3) emphasis on abstraction and modularity for reasoning. We discuss related lines (1) and (2) as well as other related work in Appendix E and emphasize reasoning and abstraction here.
@@
More recent work has shifted toward reasoning-centric uses of memory, particularly test-time learning. Think-in-Memory records intermediate reasoning steps as structured triples, storing them in a locality-sensitive hash table to enable reuse in multi-step reasoning tasks (Liu et al., 2023). Buffer of Thoughts stores problem specific reasoning templates, retrieved by embedding similarity and instantiated for new inputs (Yang et al., 2024). It categorizes and summarizes solution attempts into templated insights, which are then added to memory. In contrast, Dynamic Cheatsheet maintains a unified buffer that is adapted continuously. With each problem query, the LLM updates this memory blob by rewriting the entire buffer (Suzgun et al., 2025). Rather than retrieving specific entries, the entire cheatsheet is appended to the prompt, functioning as a persistent cache of problem-solving strategies.
@@
Our work is primarily evaluated against ARC-AGI as a benchmark that simulates complex reasoning on frontier tasks without requiring frontier-level knowledge. Various techniques and findings have been developed against this benchmark. Li et al. (2024) demonstrates the efficacy of synthetic data and the complementary approaches of (i) learning to infer the underlying transformation via program synthesis and (ii) using test-time weight adaptation to directly learn the transformation function in the neural net. Akyürek et al. (2025) further investigates the test-time adaptation method, producing various insights, including that low rank adapters are poorly suited to retaining practical experience across puzzles. Our program synthesis approach can be viewed as a memory-augmented version of Wang et al. (2024)’s hypothesis search or Qiu et al. (2024)’s hypothesis refinement with execution feedback-based retry. Other recent approaches to ARC-AGI include brain-inspired architectural innovations such as Wang et al. (2025)’s hierarchical reasoning model, achieving impressive performance through a recurrent architecture that iteratively refines its output until a halt signal is predicted, together with data augmentation techniques. Finally, the relative ease of generating new ARC-AGI tasks compared to solving them is exploited in Pourcel et al. (2025)’s self-improving system that relabels incorrect programs as the correct program for a different puzzle and trains a model on its generations this way.
@@ 3.1 Problem Statement
Input: Dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{n}$ (labels $y_{i}$ optional), External memory $M$, Operations MemRead, MemWrite, GetFeedback, Update interval $k$,
@@
Algorithm 1 Inference with Continually Updating External Memory
Figure 2: Method Diagram. Implementing a memory system requires defining (1) what is stored, (2) how memory is updated, and (3) how memory is used for new queries. The key novelty in this work is emphasizing abstraction and modularity, and the corresponding design changes. In particular, we highlight that parameterization (with higher-order functions allowed and encouraged) promotes abstraction, and typed interface definitions support modularity by showing which concepts can be combined. Since these memory entries are more abstract, they also require more inference to map against new, concrete situations–whether by aligning input against the memory format in a preprocessing query, or leveraging reasoning models to explore in a directed manner.
@@
The main requirement for our memory system is that some feedback is available at test-time (e.g. via test cases or self-reflection). This condition is necessary as retaining patterns from an incorrect or otherwise flawed trace would serve to carry mistakes forward. In other words, some signal is needed to discriminate correct from incorrect traces to ensure only productive ideas and patterns are added to memory. Incorrect traces still likely contain vital signal, but error identification/credit assignment remains an open challenge that we leave to future efforts. Various real-world tasks may satisfy this requirement, as tasks are often specified with either examples showing desired behavior or evaluation criteria defining what outcomes are positive and negative. To give some examples, code completion tools have criteria like compilation success and test pass rate, and medical diagnosis tools have evaluation criteria defined as patient outcomes (e.g., survival rates). In the case of our evaluation benchmark ARC-AGI, puzzles are explicitly presented as a set of input-output examples, which, for our purposes, can serve as automatically verifiable test cases.
@@ 3.2 Memory Format and Organization
Specifying an individual memory entry’s format involves selecting fields useful for retrieval and downstream problem-solving. The organization of memory entries on a collective level is another design surface we leave for future exploration; we use a flat collection of entries for simplicity.
Figure 3: Open-Ended (OE) vs. Program Synthesis (PS) Concept Examples. An example of concepts from puzzle 9af7a82c abstracted into each concept format. OE defers to the model, while PS imposes structure to encourage abstraction/modularity. Higher order behavior is demonstrated with the “sort objects” concept taking a Callable parameter that specifies specific variations.
@@
In early iterations of automatic concept summarization from solution traces, we observed outputs to occasionally be overly specific. Consider Figure 3’s OE example that bundles ideas of counting, drawing columns, and sorting into a single entry. This overspecification inherits the same limitations as instance-level concepts: friction when matching a composite idea to new scenarios and redundancy from composite ideas sharing components.
We designed a more structured alternative to the OE format to address these issues and more strongly enforce the ideals of abstraction and modularity. In particular, we took inspiration from software engineering and functional programming, which respectively have existing solutions in modular design for code reuse and higher-level functions for composition. From this analogy, our Program Synthesis (PS) format frames concepts as types, structures, and routines. The primary feature is that each of these concepts is parameterized. Parameterization allows similar concepts to be represented compactly with variations abstracted into parameters. Moreover, type annotations for inputs (parameters) and outputs (return types) encourage modularity and composition: the typed interfaces suggest what ideas fit together and how. Finally, by allowing parameters to be routines themselves (introducing higher-order functions), we encourage further generalization by enabling abstraction over specialized logic and routines rather than just values. Higher-level patterns can be recognized across instances and stored with routine arguments defining the different lower-level logic. A subtle benefit of this more structured formulation is that memory representation can be easily compressed by omitting certain fields (see subsection A.1 for a complete list of PS concept fields).
@@ 3.3 Memory Write: Concept Abstraction
Memory’s primary purpose is to persist discoveries and reflections from prior experiences. Converting experiences to memory updates is then the most crucial operation for achieving persistence.
@@
Abstracting suggestions from a solution trace is straightforward–simply query a model to reflect on the solution trace and summarize specific general ideas that may be reused for future puzzles. To synthesize specific conditions or situational cues to pair with suggestions, ideally, we can directly refer to the original problem-solving process as a series of deductions that derive new facts or explore ideas from existing facts, and link each suggestion (idea) with the prior observations that inferred this suggestion. However, intermediate inferences may be implicit or difficult to extract from a trace, or, in the case of specific commercial models, completely hidden. To address this challenge, we generate a post-hoc derivation: an interleaved sequence of observations and thoughts/reasoning steps constructed from the input-output of the final solution. These reconstructed traces are then used to extract situation-suggestion pairs, forming structured memory entries for guiding future analyses. This minimal formulation does not consider the existing contents of memory while adding new entries for simplicity and scaling reasons. It defers handling redundancy and consolidating ideas to the concept selection and problem-solving phases, respectively.
@@
We find that directly converting solutions into routines often leads to minor implementation details being recorded as memories, as opposed to our intended abstract concepts. We mitigate this by preprocessing solutions into pseudocode to prioritize higher-level operations over low-level implementation details. The main abstraction phase proceeds from the generated pseudocode, recording new concepts and revising existing concepts along with their various fields.
@@ 3.4 Memory Read: Concept Selection
Although state-of-the-art models support longer context windows, their maximum length still limits how much information or how many memory entries can be included in a single query. Even without this hard limit, including all available memories in context is still undesirable: it can distract the model with irrelevant details and flood it with too many unlikely hypotheses to consider. A more effective strategy is to introduce a selection mechanism only to include the most relevant subset of memory entries at problem-solving time. This allows the memory store to grow continually without overwhelming the model’s context window.
@@
OE-format memory entries were explicitly designed to support memory-based selection. The situation clause acts as a semantic hook to identify relevance. The key idea for OE Selection is a preprocessing step that leverages model reasoning capability to parse the abstract domain input into problem descriptions at different levels of abstraction to facilitate matching against our abstract concepts. Since ARC-AGI deals with spatial reasoning, we leverage a vision language model for this preprocessing step. We caption each puzzle using a structured prompt that separates concrete observations from speculative transformations. This converts spatially rich input into a natural language format suitable for matching against stored situations. We then query a model for the top-$k$ most relevant entries using the generated description. We explored scoring, thresholding, and cumulative similarity approaches (as in top-$p$ sampling). Observing comparable results, we selected top-$k$ cutoff for simplicity. We consider this a light-weight approach to converting abstract domain inputs into more understandable and memory-aligned representations.
@@
Standard embedding approaches use a single forward pass per embedding, representing a System 1-style fast thinking/intuition-driven understanding from the model. Intuitions may not be sufficient for difficult/frontier tasks and domains. Moreover, the connection between a higher-level concept and a specific problem instance may not be immediately clear because of the abstraction.
To address these challenges, we propose reasoning-based selection–using System 2-style thinking to deliberately think and explore. In contrast to the more straightforward playbooks curated in the OE memory format, the explicitly abstract PS concepts are a set of puzzle pieces with notes on identifying if a piece is relevant (relevance cues) and on how various pieces fit together (type annotations). To select from this memory, we propose leveraging model-driven exploration instantiated by recent models’ long-form reasoning with backtracking. PS Selection instructs a reasoning model to systematically explore the problem: first identify initial concepts using relevance cue annotations, then attempt to “fill in the details” by determining values or routines to populate these initial concepts’ parameters using type annotations to identify which other concepts should be investigated.